OpenStudy (anonymous):

what methods are useful to create the equation of a line? Demonstrate by writing an equation in standard form parallel to y=3x+2 through the point (5,-6)

OpenStudy (johnweldon1993):

Well we know parallel lines have the same slope...so we can start with that.. y = 3x + 2 This is slope intercept form y = mx + b where 'm' is the slope of the line... So there....'m' = 3 so the slope is 3...just like the line we are looking for... We need our line to pass through (5 , -6) (x , y) We can plug in the 'm' we have...and also the point we have into the point slope formula \[\large y - y_0 = m(x - x_0)\] After plugging in we have \[\large y + 6 = 3(x - 5)\] \[\large y = 3x - 21\] This would be our line...and to put this in standard form we just subtract the 3x from both sides to get this to look like \[\large Ax + By = C\] \[\large -3x + y = -21\] and that is our line

OpenStudy (anonymous):

one more question. :/ sorry guys!! solve f(x)=x^2-4x+3 for f(-2)

OpenStudy (anonymous):

and thank you sooooo much!!! :D

OpenStudy (johnweldon1993):

if \[\large f(x) = x^2 - 4x + 3\] f(-2) would be that same equation...except wherever you see an 'x' you replace it with (-2) so \[\large f(-2) = (-2)^2 - 4(-2) + 3\] \[\large f(-2) = 4 - (-8) + 3\] \[\large f(-2) = 4 + 8 + 3 = \space ?\]

OpenStudy (anonymous):

eehhmmmmm? 0.o need I say that im LITERALLY brain dead with math.. :/ sorry.

OpenStudy (johnweldon1993):

Lol don't worry about it :) Alright so lets start over \[\large f(x) = x^2 - 4x + 3\] this is our original equation right?

OpenStudy (johnweldon1993):

And we need to find \[\large f(-2)\]

OpenStudy (johnweldon1993):

See how it went from \[\large f(x) \rightarrow f(-2)\] the only thing that changed...is where 'x' once was...it is now -2 So in order to solve our equation...that's all WE need to do too...replace every 'x' we see in the equation...and replace it with -2

OpenStudy (johnweldon1993):

\[\large f(x) = x^2 - 4x + 3\] now becomes \[\large f(-2) = (-2)^2 - 4(-2) + 3\] All that happened was I replaced \(\large x^2\) with \(\large (-2)^2\) and \(\large -4x\) with \(\large -4(-2)\)

OpenStudy (johnweldon1993):

So now we have \[\large f(-2) = (-2)^2 - 4(-2) + 3\] \[\large f(-2) = 4 - (-8) + 3\] \[\large f(-2) = 4 + 8 + 3\] \[\large f(-2) = 15\]

OpenStudy (anonymous):

omg thats alot easier after you break it up. :3 thank you!!

OpenStudy (anonymous):

what processes are needed to solve for functions and thier inverses?

OpenStudy (anonymous):

needs to be done in under a min >.<

OpenStudy (johnweldon1993):

To solve for a functions inverse...you need to 1) switch the 'x' and the 'y' in the original function 2) solve for 'y' again